# Assessment of the Earth orientation parameter accuracy from concurrent VLBI observations

### v1

@@ 4.1 General geodetic data analysis
The general VLBI data analysis procedure is presented in full detail in Petrov and Kovalev [22]. To avoid unnecessary repetition, we present here only highlights. We applied 3D displacements due to ocean tidal, ocean non-tidal, atmospheric pressure, and land water storage loadings provided by the International Mass Loading Service [20] and computed from the outputs of NASA Global Modeling Office numerical weather models. We applied in data reduction a priori slant path delays computed by a direct integration of equations of wave propagation through the heterogeneous atmosphere [19] using the output of NASA numerical weather model GEOS-IT which is a successor of the older model GEOS-FPIT [26, 14, 27].
@@
We estimated the following global parameters: 1) positions of all the stations at the reference epoch 2000.01.01; 2) linear velocities of all the stations; 3) antenna axis offsets of most of the stations; 4) sine and cosine components of harmonic site position variations of 69 stations with a long history of observations at diurnal, semi-diurnal, annual, and semi-annual frequencies; 5) B-spline coefficients that model the non-linear motion of 29 stations that included sudden co-seismic position changes, smooth post-seismic relaxation, non-linear change of the antenna tilt, non-linear uplift due to glaciers melting, non-linear local motions, and discontinuities due to station repair; 6) coefficients of harmonic variations of the EOP vector at 890 frequencies including nutations; and 7) positions of all the sources with at least three usable observations.
@@
When processing 24-hr experiments, we estimated Euler angles at the specified epoch and their time derivatives as local parameters. We also estimated clock breaks and baseline-dependent clock for some experiments with clock function discontinuities or baseline-dependent biases. In addition, we estimated B-spline coefficients of the 1st degree for residual atmospheric path delay in the zenith direction, tilts of the symmetry axis of the refractivity field, also known as atmospheric gradients, and clock functions for all the stations, except the one taken as a reference, as segmented parameters. Time spans of B-splines were 20 minutes for atmospheric zenith path delay, 6 hours for tilts, and 60 minutes for clock function.
@@
When processing 1-hr experiments, we estimated Euler angle $E_{3}$ as a local parameter, clock function and residual atmospheric path delay as segmented parameters. Time spans of B-splines were 10 minutes for atmospheric zenith path delay and 20 minutes for clock function.
@@
We also imposed constraints on estimated parameters. An exhaustive description of constraints can be found in Petrov and Kovalev [22]. Here we stress that Earth Orientation Parameters were unconstrained, and only minimum constraints were imposed on the station positions and their derivatives. The motivation for our choice of running global solutions with estimating all parameters, including EOP, in a single least squares run, rather than fixing some parameters, is that this automatically imposed consistency of the estimated station positions and EOP. It was reported in the past that fixing nutation, polar motion, or station positions to some a priori values that are considered “good” resulted in a measurable degradation of statistics because of inconsistency that a prior models [16, 6]. In addition, by solving for station positions and source coordinates, we effectively incorporate them in a solution as quantities that have uncertainties described by the full covariance matrix, while fixing them is equivalent to treating them as known exactly.
@@
The $\chi^{2}$ per the number of degrees of freedom statistics shows that group delay uncertainties computed by the fringe fitting algorithm, $\sigma_{a}$, never match post-fit residuals. Using weights $w=1/\sigma_{a}$ results in very significant biases of estimates of standard variations of adjusted parameters. To alleviate this problem, we used weights in the form $w=1/\sqrt{\sigma_{a}^{2}+v^{2}}$, where $v$ is an extra variance computed for each experiment and for each baseline in such a way that the ratio of the weighted sum of postfit residuals to their mathematical expectation is close to unity following the approach first described in Ryan et al. [30]. The mathematical apparatus of this technique is presented in Petrov and Kovalev [22]. This additional variance $v$ turned out close to the weighted root mean square (wrms) of postfit residuals for many experiments.
@@ 5.1 Results of s22 campaign
We computed biases and rms of differences in $E_{3}$ angles from each 22-hr observing session and from all the data combined. Results are presented in Table 2. Errors of both r1r4 and s22 campaigns contribute to the rms of differences. We corrected these rms by subtracting in quadrature the contribution to these differences from 24-hr programs that was set to 0.53 nrad for all the experiments. We discuss how we obtained this contribution in subsection 5.3. We call these statistics “corrected”.
Even a cursory analysis reveals a noticeable difference between experiments ran in summer months May, June, July, August, September and winter months November, December, January, February, and April. According to the classical F-test, the ratio of standard deviations, 1.85 is statistically significant at a $10^{-9}$ confidence level for a sample of this size.
Table 2: Estimates of biases and rms of differences in $E_{3}$ angles determined from 1-hr s22 experiments with respect to concurrent multi-baseline 24-hr experiment. The rms of the differences was corrected for the contribution of errors from 24-hr experiments.
Session name Session code bias rms # exp
@@
Winter 0.25 0.78 127
Summer 0.01 1.44 88
All 0.17 1.04 215
Figure 4 illustrates the differences in $E_{3}$ angles from two observing sessions with the smallest and largest standard deviations. The differences do not exhibit a noticeable systematic pattern. The biases in $E_{3}$ angles do not exceed 60% of a standard deviation for any observing session and they are at a level of 1/6 of the standard deviation of the whole dataset.
We computed the rms of the differences for experiments scheduled according to strategy A and strategy B separately. We got 1.04 and 1.05 nrad respectively, which is totally insignificant. The differences remained insignificant when winter only or summer only experiments were used.
@@
5.2 Results from regular 1-hr observing campaign at KkWz
The goal of 1-hr campaigns is to provide $E_{3}$ angle estimates with a short latency. For instance, the goal of IVS-INT-00 campaign is to have a latency of 36 hours, and this explains why these campaigns are called intensive. To meet these requirements, the experiment duration is reduced to one hour and the number of stations is reduced to two. Occasionally, a third and fourth station is added. A significant number of these experiments overlap with r1r4 24-hr observing sessions. We set EOP reference epochs for corresponding r1r4 experiments to the middle epoch of 1-hr experiments and derive $E_{3}$ angles at exactly the same epochs from 1-hr and 24-hr experiments. We computed statistics of the differences and applied an iterative procedure starting with the data point with the greatest $|x|/$rms ratio and consecutively eliminating such points and re-computing the rms till no data points with this ratio greater than 3 remained. The number of outliers does not exceed 2%. The statistics of the differences corrected for outliers, as well as the mean values of the corresponding formal errors are shown in Table 3.
@@ 5.2 Results from regular 1-hr observing campaign at KkWz
K2Ws 409 0.29 1.30 0.59
MgWs 29 -0.17 1.22 0.62
MgWs 215 0.17 1.04 0.66
TsWz 107 0.21 1.45 0.99
IsWz 65 0.74 1.43 0.86
MkWz 115 0.04 1.21 0.72
HnMk 58 0.75 1.60 1.36
KkWz 2905 -0.22 1.50 1.26
@@
We show in Figure 6 the differences in $E_{3}$ angles from the INT-INT-1 campaign that has been running for over 25 years. We do not see a significant systematic pattern in the plot of residuals. Then we computed the rms over six summer months May, June, July, August, September, October and over six winter moths. The rms of differences are 1.57 and 1.39 nrad respectively. Since the number of winter and summer concurrent experiments is over 1400, the F-test shows that this difference is significant at a $10^{-5}$ level. Winter IVS-INT-1 experiments are more accurate than summer experiments. Summer experiments on average add in quadrature 0.78 nrad variance.
@@
We performed a similar summer/winter test for VGOS-INT-A campaign at baseline K2Ws. Both ends of baseline K2Ws and KKWz are located within less than 100 meters of each other. The corrected rms of the differences are 1.11 nrad for winter and 1.42 nrad for summer. According to F-test these differences are significant at a $6\cdot 10^{-4}$ level. Summer experiments on average add in quadrature 0.88 nrad variance.
@@ 5.3 Results from regular concurrent 24-hr VLBI campaigns
$E_{1}$, $E_{2}$, and $\dot{E}_{3}$ are regularly estimated during analysis of GNSS data by a number of groups. This prompted us to use the three corner hat method [8] for evaluation of uncertainties of these parameters from GNSS, vo, and r1r4 campaigns. Indeed, when three time series are uncorrelated, we can derive the variances of each series from the variances of their differences. We took daily time series of polar motion and lengths of day from the IGS final series, converted them to $E_{1}$, $E_{2}$, and $\dot{E}_{3}$, interpolated to the epoch of observations, computed the rms of the differences, and then derived standard deviations of each three series. Results are presented in Table 5. We are aware that the three corner hat method is susceptible to the presence of weak correlation and systematic errors. We computed a smoothing spline at 9 equally distributed knots to fit the difference between the series and removed a small systematic pattern, but this did not change the results noticeably.
We see that the accuracy of polar motion and $\dot{E}_{3}$ from GNSS has improved by a factor of 2.5 over 25 years since 1997–2000.5. Our accuracy estimate, 0.20 nrad for $E_{1}$ and $E_{2}$, is 50% higher than the estimate of Ray et al. [25] (0.12–0.15 nrad). The EOP accuracy from VLBI also improved, but the improvement is not that strong: 28 to 68%. The results of our comparison shows that accuracy of $E_{1}$ and $E_{2}$ from r1r4 sessions is better than both accuracy from old ca/na campaigns and from vo campaign. According to the F-test, this accuracy difference is significant at a level of 0.002. The difference in uncertainties of $\dot{E}_{3}$ from these campaigns is significant at the 0.01 level. Our tests confirm that accuracy of $E_{1}$ and $E_{2}$ from GNSS was on par with VLBI in late 1990s, but it improved by a factor of 2–3 by the 2020s.
@@ 5.4 Results from 1-hr blocks of concurrent 24-hr VLBI campaigns
$f(t)=\left\{\begin{array}[]{l}t^{\beta_{1}}\enskip{\rm if}\enskip t>b\vspace{1ex}\\ t^{\beta_{2}}\enskip{\rm if}\enskip t\leq b\end{array}\right.,$ (10)
@@
Figure 8: Rms of errors in estimates of $E_{1}$ (left) and $E_{3}$ (right) angles as a function of experiment duration from vo/r1 comparison (green dots). The middle blue line shows the best fit of the broken power laws, while the bottom black line shows the power law with $\beta=-0.50$, representing the expectation based on uncorrelated observations.
@@ 6.1 Outcomes of different scheduling strategies
Within the s22 experiments, we explicitly tested an observing strategy that we expected would provide a more precise evaluation of the atmospheric path delay from the observations themselves, because we included observations of sources at high and low elevation angles in a rapid sequence over short periods of time. We generated simulated datasets from schedules and added the noise based on the mathematical model of atmospheric turbulence and its code developed by Pany et al. [17]. We ran a battery of least square solutions of the simulated datasets with a diagonal weight matrix, generated time series of simulated $E_{3}$ estimates and computed their rms: 1.24 and 1.21 nrad for scheduling strategies A and B respectively.
@@
10 min segments 0.22 1.04 0.74 25.61
60 min segments 0.19 1.04 0.69 29.89
no estimation 3.35 1.46 0.66 56.31
@@ 6.2 Seasonal accuracy variability
Putting together these two findings, we conclude that the dominant error source in EOP estimates is due to the residual contribution of the atmospheric path delay that was not captured by the estimation of zenith path delay and tilts of the refractivity field symmetry axis. Formal uncertainties of group delays in quad-band experiments are at a level of 3–5 ps all year round, while the scatter of postfit residuals varies in a range of 15–25 ps. There should be some source of these errors that is greater in summer than in winter. Path delay is greater in summer. If the atmospheric path delay is entirely captured by the estimation process, an increase in wet path delay would not affect residuals. When estimating zenith path delay and tilts captures only some share of the total signal, postfit residuals will depend on the total path delay. The unmodeled part of path delay will affect estimates of other parameters, such as EOP. The evidence we presented supports this explanation.
@@ 6.3 EOP accuracy as a function of an observing session duration
VLBI data that were used for deriving angles $E_{3}(t_{1})$ and $E_{3}(t_{2})$ from 0.5 hr blocks at $t_{1}$ and $t_{2}$ epochs are independent. The thermal noise in these data is independent and uncorrelated. If the only contribution to the noise were the thermal noise in receivers, then rms of the difference in $E_{3}$ estimates, $\sigma_{0}$, would have been $\sqrt{2}\,\sigma_{0}$ of the rms of individual $E_{3}$ estimates from 0.5 hr segments, i.e. 2 $\sqrt{2}\>\cdot\>1.42=2.01$ nrad. But we got $\sigma=1.47$ nrad, which is substantially less. We can compute correlation $\rho$ between estimates of $E_{3}$ angles derived from 0.5 hr date segments using a simple expression

### v2

@@ 4.1 General geodetic data analysis
The general VLBI data analysis procedure is presented in full detail in r:rfc1. To avoid unnecessary repetition, we present here only highlights. We applied 3D displacements due to ocean tidal, ocean non-tidal, atmospheric pressure, and land water storage loadings provided by the International Mass Loading Service [r:malo15] and computed from the outputs of NASA Global Modeling Office numerical weather models. We applied in data reduction a priori slant path delays computed by a direct integration of equations of wave propagation through the heterogeneous atmosphere [r:padel] using the output of NASA numerical weather model GEOS-IT which is a successor of the older model GEOS-FPIT [r:reichle11, r:geos12, r:geos18].
@@
We estimated the following global parameters: 1) positions of all the stations at the reference epoch 2000.01.01; 2) linear velocities of all the stations; 3) antenna axis offsets of most of the stations; 4) sine and cosine components of harmonic site position variations of 69 stations with a long history of observations at diurnal, semi-diurnal, annual, and semi-annual frequencies; 5) B-spline coefficients that model the non-linear motion of 29 stations that included sudden co-seismic position changes, smooth post-seismic relaxation, non-linear change of the antenna tilt, non-linear uplift due to glaciers melting, non-linear local motions, and discontinuities due to station repair; 6) coefficients of harmonic variations of the EOP vector $P_{c},P_{s},S_{c},S_{s},R_{c},R_{s}$ at 890 frequencies including nutations; and 7) positions of all the sources with at least three usable observations.
When processing 24-hr experiments, we estimated Euler angles $E_{1s},E_{2s},E_{3s}$ at the specified epoch $t_{\rm ref}$ and their time derivatives as local parameters:
[…]
as well as global parameters mentioned above. Here we added symbol $(s)$ to emphasize that these parameters do not depend on time, but on the specific observing session $s$. Relationship between estimated parameters $E_{1}(s),E_{2}(s),E_{3}(s)$ and parameters $X_{p},Y_{p},$ UT1-UTC used by IERS are given in appendix B.
@@
We also estimated clock breaks and baseline-dependent clock for some experiments with clock function discontinuities or baseline-dependent biases. In addition, we estimated B-spline coefficients of the 1st degree for residual atmospheric path delay in the zenith direction, tilts of the symmetry axis of the refractivity field, also known as atmospheric gradients, and clock functions for all the stations, except the one taken as a reference, as segmented parameters. Time spans of B-splines were 20 minutes for atmospheric zenith path delay, 6 hours for tilts, and 60 minutes for clock function.
@@
When processing 1-hr experiments, in addition to global parameters, we estimated Euler angle $E_{3}$ as a local parameter, clock function and residual atmospheric path delay as segmented parameters:
[…]
as well as global parameters, and kept $E_{1},E_{2}$ fixed. Time spans of B-splines were 10 minutes for atmospheric zenith path delay and 20 minutes for clock function. We used this relatively short time spans because we identify the presence of the atmospheric signal in residuals of some experiments when we used longer time spans.
@@
We also imposed constraints on estimated parameters. An exhaustive description of constraints can be found in r:rfc1. Here we stress that Earth Orientation Parameters were unconstrained, and only minimum constraints were imposed on the station positions and their derivatives. The motivation for our choice of running global solutions with estimating all parameters, including EOP, in a single least squares run, rather than fixing some parameters, is that this automatically imposed consistency of the estimated station positions and EOP. Estimation of nutation daily offset from processing 1-hr experiments is not feasible. It was reported in the past that fixing nutation, polar motion, or station positions to some a priori values that are considered “good” resulted in a measurable degradation of statistics because of inconsistency that a prior models [r:not08, r:dieck23]. Our approach of estimating harmonic EOP variations as global parameters overcomes this limitation. In addition, by solving for station positions and source coordinates, we effectively incorporate them in a solution as quantities that have uncertainties described by the full covariance matrix, while fixing them is equivalent to treating them as known exactly.
@@
The $\chi^{2}$ per the number of degrees of freedom statistics shows that group delay uncertainties computed by the fringe fitting algorithm, $\sigma_{a}$, never match post-fit residuals. Using weights $w=1/\sigma_{a}$ results in very significant biases of estimates of   variances of adjusted parameters. To alleviate this problem, we used weights in the form $w=1/\sqrt{\sigma_{a}^{2}+v^{2}}$, where $v$ is an extra variance computed for each experiment and for each baseline in such a way that the ratio of the weighted sum of postfit residuals to their mathematical expectation is close to unity following the approach first described in r:ryan91. The mathematical apparatus of this technique is presented in r:rfc1. This additional variance $v$ turned out close to the weighted root mean square (wrms) of postfit residuals for many experiments.
@@ 5.1 Results of s22 campaign
We computed biases and rms of differences in $E_{3}$ angles from each 22-hr observing session and from all the data combined. Results are presented in Table 2. Errors of both r1r4 and s22 campaigns contribute to the rms of differences. We corrected these rms by subtracting in quadrature the contribution to these differences from 24-hr programs that was set to   0.40 nrad for all the experiments. We discuss how we obtained this contribution in subsection 5.3. We call these statistics “corrected”.
Even a cursory analysis reveals a noticeable difference between experiments ran in summer months May, June, July, August, September and winter months November, December, January, February, and April. According to the classical F-test, the ratio of   rms, 1.85 is statistically significant at a $10^{-9}$ confidence level for a sample of this size.
Table 2: Estimates of biases and   stdof differences in $E_{3}$ angles determined from 1-hr s22 experiments with respect to concurrent multi-baseline 24-hr experiment. The   std of the differences was corrected for the contribution of errors from 24-hr experiments.
Session name Session code bias   std # exp
@@
Winter 0.29 0.65 127
Summer 0.07 1.42 88
All 0.22 0.97 215
Figure 3 illustrates the differences in $E_{3}$ angles from two observing sessions with the smallest and largest   rms. The differences do not exhibit a noticeable systematic pattern. The biases in $E_{3}$ angles do not exceed 40% of a   rms for any observing session and they are at a level of 1/5 of the   rms of the whole dataset.
We computed the rms of the differences for experiments scheduled according to strategy A and strategy B separately. We got 0.96 and 0.97 nrad respectively, which is totally insignificant. The differences remained insignificant when winter only or summer only experiments were used.
@@
5.2 Results from regular 1-hr observing campaigns
The goal of 1-hr campaigns is to provide $E_{3}$ angle estimates with a short latency. For instance, the goal of IVS-INT-00 campaign is to have a latency of 36 hours, and this explains why these campaigns are called intensive. To meet these requirements, the experiment duration is reduced to one hour and the number of stations is reduced to two. Occasionally, a third and fourth station is added. A significant number of these experiments overlap with r1r4 24-hr observing sessions. We set EOP reference epochs for corresponding r1r4 experiments to the middle epoch of 1-hr experiments and derive $E_{3}$ angles at exactly the same epochs from 1-hr and 24-hr experiments. We computed statistics of the differences and applied an iterative procedure starting with the data point with the greatest |x|/|x|/rms ratio and consecutively eliminating such points and re-computing the rms till no data points with this ratio greater than 3 remained. The number of outliers does not exceed 2%. The statistics of the differences corrected for outliers, as well as the mean values of the corresponding formal errors are shown in Table 3.
@@ 5.2 Results from regular 1-hr observing campaigns
K2Ws 409 0.29   1.40 0.59
MgWs 29 -0.17   1.33 0.62
MgWs 215 0.17   0.97 0.66
TsWz 107 0.21   1.54 0.99
IsWz 65 0.74   1.53 0.86
MkWz 115 0.04   1.32 0.72
HnMk 58 0.75   1.68 1.36
KkWz 2905 -0.22   1.59 1.26
@@
We show in Figure 5 the differences in $E_{3}$ angles from the   IVS-INT-1 campaign that has been running for over 25 years. We do not see a significant systematic pattern in the plot of residuals. Then we computed the   std over six summer months May, June, July, August, September, October and over six winter moths. The rms of differences are   1.66 and 1.49 nrad respectively. Since the number of winter and summer concurrent experiments is over 1400, the F-test shows that this difference is significant at a $10^{-5}$ level. Winter IVS-INT-1 experiments are more accurate than summer experiments. Summer experiments on average add in quadrature 0.73 nrad variance.
@@
We performed a similar summer/winter test for VGOS-INT-A campaign at baseline K2Ws. Both ends of baseline K2Ws and KKWz are located within less than 100 meters of each other. The corrected rms of the differences are   1.23 nrad for winter and 1.52 nrad for summer. According to F-test these differences are significant at a $6\cdot 10^{-4}$ level. Summer experiments on average add in quadrature 0.89 nrad variance with respect to winter experiments.
@@ 5.3 Results from regular concurrent 24-hr VLBI campaigns
$E_{1}$, $E_{2}$, and $\dot{E}_{3}$ are regularly estimated during analysis of GNSS data by a number of groups. This prompted us to use the three corner hat method [r:tch] for evaluation of uncertainties of these parameters from GNSS, vo, and r1r4 campaigns. Indeed, when three time series are uncorrelated, we can derive the variances of each series from the variances of their differences. We took daily time series of polar motion and lengths of day from the IGS final series, converted them to $E_{1}$, $E_{2}$, and $\dot{E}_{3}$, interpolated to the epoch of observations, computed   the bias and rms of the differences, and then derived standard deviations of each three series. Results are presented in Table 5. We are aware that the three corner hat method is susceptible to the presence of weak correlation and systematic errors. We computed a smoothing spline at 9 equally distributed knots to fit the difference between the series and removed a small systematic pattern, but this did not change the results noticeably.
We see that the accuracy of polar motion and $\dot{E}_{3}$ from GNSS has improved by a factor of 2.5 over 25 years since 1997–2000.5. Our accuracy estimate, 0.20 nrad for $E_{1}$ and $E_{2}$, is 50% higher than the estimate of r:ray17 (0.12–0.15 nrad). The EOP accuracy from VLBI also improved, but the improvement is not that strong: 28 to 68%. The results of our comparison shows that accuracy of $E_{1}$ and $E_{2}$ from r1r4 sessions is better than both accuracy from old ca/na campaigns and from vo campaign. According to the F-test, this accuracy difference is significant at a level of 0.002. The difference in uncertainties of $\dot{E}_{3}$ from these campaigns is significant at the 0.01 level. Our tests confirm that accuracy of $E_{1}$ and $E_{2}$ from GNSS was on par with VLBI in late 1990s, but it improved by a factor of 2–3 by the 2020s.
@@ 5.4 Results from 1-hr blocks of concurrent 24-hr VLBI campaigns
$f(t)=\left\{\begin{array}[]{l}t^{\beta_{1}}\enskip{\rm if}\enskip t>b\vskip 4.30554pt\\ t^{\beta_{2}}\enskip{\rm if}\enskip t\leq b\end{array}\right$., (22)
@@
Figure 7: The rms of errors in estimates of $E_{1}$ (left) and $E_{3}$ (right) angles as a function of experiment duration from vo/r1 comparison (green dots). The middle blue line shows the best fit of the broken power laws, while the bottom black line shows the power law with $\beta=-0.50$, representing the expectation based on uncorrelated observations.
@@ 6.1 Outcomes of different scheduling strategies
Within the s22 experiments, we explicitly tested an observing strategy that we expected would provide a more precise evaluation of the atmospheric path delay from the observations themselves, because we included observations of sources at high and low elevation angles in a rapid sequence over short periods of time. We generated simulated datasets from schedules and added the noise based on the mathematical model of atmospheric turbulence and its code developed by r:pany11. We ran a battery of least square solutions of the simulated datasets with a diagonal weight matrix, generated time series of simulated $E_{3}$ estimates and computed their rms: 1.24 and 1.21 nrad for scheduling strategies A and B respectively.
@@
10 min segments 0.22   0.97 0.74 25.61
60 min segments 0.19   0.97 0.69 29.89
no estimation 3.35   1.49 0.66 56.31
@@ 6.2 Seasonal accuracy variability
Putting together these two findings, we conclude that the dominant error source in EOP estimates is due to the residual contribution of the atmospheric path delay that was not captured by the estimation of zenith path delay and tilts of the refractivity field symmetry axis. Formal uncertainties of group delays in quad-band experiments are at a level of 3–5 ps all year round, while the scatter of postfit residuals varies in a range of 15–25 ps. There should be some source of these errors that is greater in summer than in winter. Path delay is greater in summer. If the atmospheric path delay is entirely captured by the estimation process, an increase in wet path delay would not affect residuals. When estimating zenith path delay and tilts captures only   a certain share of the total signal that depends on schedule, postfit residuals and the impact of the unmodeled atmospheric path delay will depend on the total path delay. The unmodeled   share of path delay will affect estimates of other parameters, such as EOP. The evidence we presented supports this explanation.
@@ 6.3 EOP accuracy as a function of an observing session duration
VLBI data that were used for deriving angles $E_{3}(t_{1})$ and $E_{3}(t_{2})$ from 0.5 hr blocks at $t_{1}$ and $t_{2}$ epochs are independent. The thermal noise in these data is independent and uncorrelated. If the only contribution to the noise were the thermal noise in receivers, then rms of the difference in $E_{3}$ estimates, $\sigma_{0}$, would have been $\sqrt{2}\,\sigma_{0}$ of the rms of individual $E_{3}$ estimates from 0.5 hr segments, i.e. $\sqrt{2}\>\cdot\>1.42=2.01$ nrad. But we got $\sigma=1.47$ nrad, which is substantially less. We can compute correlation $\rho$ between estimates of $E_{3}$ angles derived from 0.5 hr date segments using a simple expression
@@
6.4 Impact of source structure
r:and18 processed a 15 day VLBI observing campaign CONT14, analyzed group delay misclosures over each triplet of stations, and found that among 72 observed sources, 6 had small rms of group delay misclosures, less than 10 ps. The rms of postfit residuals over this group of sources was 19.2 ps versus 25.2 ps over all the sources. Based on this finding, the authors made a conclusion that source-structure-related errors in geodetic VLBI are about as large as all other error sources combined. However, if two processes have a similar impact on residuals, this does not necessarily mean that their impact on EOP estimates is the same. Here we make a rough estimate of the source stucture contriubution on EOP determination.
Since we do not have an adequate model for source structure, we can evaluate the impact of source structure contribution only indirectly. Simulations of r:pla16 and our prior work for determining source position using available source images [r:gaia3] demonstrated that source structure affects source positions mainly along jet directions.
The contribution of structure contribution causes a jitter in source position estimates. That process can be characterized by the first two moments $\mu$ and $\sigma^{2}$. We should note that the first moment $\mu$ is absorbed in source position estimates and therefore, does not impact EOP estimates.
In r:radf we estimated the rms of source position differences along the jet direction derived from observations at 24 GHz with respect to positions derived from dual-band 8.4/2.3 GHz observations: 0.60 nrad. We consider this as an estimate of the first moment of the source position jitter $\mu$ for the entire population of observed sources. We attribute these differences to the contribution of source structure. We consider with some reservations this estimate as a measure of the impact of source structure on source positions on average. Although the differences in source positions from two bands are not entirely from source source structure, which would cause an overestimation of the impact, and source structure at two bands may be in part common, which would cause an underestimation of the impact, this is the best estimate of the impact based on evidence that we have now.
Only a portion of source structure contribution propagates to source position estimate, the remaining contribution propagates to the residuals. But we assume here that only that portion that affects source position may affect EOP estimates.
If source position offset along the jet direction were the Gaussian process, the first two moments $\mu$ and $\sigma^{2}$ had been independent. But that would mean that for a given source position change due to source stucture would occur in both directions along the jet with an equal probabiulity, i.e. jets would randomly change sign of their direction. Observations showed [see, for instance, r:mojave_xvii] that this never happens. For one-side distributions, moments are not independent. Since we do not know what is the actual distribution of source position variations due to source structure change, following the Occam principle, we consider two simplest distributions, the half-normal and Rayleigh, which correspond to distributions of $|X|$ and $\sqrt{X^{2}+Y^{2}}$, where $X$, $Y$ are normally distributed zero mean variables. The ratio of moments $\sigma/\mu$ for these distributions is $\sqrt{\frac{\pi-2}{2}}$ and $\sqrt{\frac{4-\pi}{\pi}}$, respectively, or 0.60 and 0.76. This estimates allow us to derive the upper limit of the variance in source positions due to the unaccounted source structure contribution at 8.4 GHz: 0.4 nrad.
To quantify the magnitude of the source structure contribution on EOP, we performed a simulation. We added the random noise to source positions with the zero mean and rms $1$ nrad along random uniformly distributed directions. The seed of the random noise was reset for each observing session. We found that a jitter in source position with a given rms $j$ affects the vector EOP estimates from r1r4 and vo networks as $(0.32,0.23,0.37)\cdot j$. These coefficients depend on the network and VLBI schedule. This allows us to provide a rough estimate of the impact of source structure on EOP: $\sim\!0.1$ nrad.
Considering that the EOP accuracy from 24-hr experiments is 0.4–0.6 nrad, we conclude that the impact of source unmodeled structure is not measurable.